# The late Miocene 10Be anomaly and the possibility of a supernova

### v1

@@
Recent measurements of cosmogenic 10Be in deep-ocean ferromanganese crusts from the Central and Northern Pacific have revealed an anomalous concentration between 11.5 and 9.0 Myr ago, peaking at 10.1 Myr. One possible explanation is a nearby supernova (SN) event. Motivated by this and by the proximity of the Solar System to the Orion star-forming region during that period, we estimate the probability that at least one SN occurred between the onset and peak of the anomaly. Using an open cluster catalog based on Gaia DR3, we trace back the orbits of 2725 clusters and the Sun over the past 20 Myr and compute the expected number of SN events. We find 19 clusters with a probability greater than 1% each of producing at least one SN within 100 pc of the Sun in the time interval 11.5–10.1 Myr ago. The total cumulative probability exceeds zero at 35 pc from the Sun and increases rapidly with distance, reaching 68% near 100 pc. Two young clusters dominate the SN probability: ASCC 20 contributes most within 70 pc, while OCSN 61 becomes more significant beyond that distance. Our results support the plausibility of a SN origin for the 10Be anomaly and highlight the importance of additional 10Be records from independent terrestrial archives to determine whether the anomaly is of astrophysical or terrestrial origin.
@@
The Solar System orbits the center of the Milky Way, together with billions of other stars and vast reservoirs of interstellar gas. Characterizing past environments crossed by the Sun, including possible encounters with large-scale Galactic structures or nearby supernovae (SNe), helps to find possible connections between Galactic environments and Earth’s geological record, fostering interdisciplinary research (see e.g., Fuchs et al., 2006; Breitschwerdt et al., 2016; Koll et al., 2019; Wallner et al., 2021; Miller & Fields, 2022; Opher et al., 2024; Maconi et al., 2025; Zucker et al., 2025).
@@
From an astronomical perspective, the European Space Agency’s Gaia mission (Gaia Collaboration et al., 2016) has enabled significant advancements, revolutionizing our understanding of the local Galactic environment. For example, Gaia data have been used to unveil the 3D structure of the solar neighborhood (see e.g., Leike et al., 2020; Vergely et al., 2022; Edenhofer et al., 2024), compile new molecular cloud catalogs (see e.g., Zucker et al., 2019; Cahlon et al., 2024), and significantly expand the open cluster census (see e.g., Castro-Ginard et al., 2018; Cantat-Gaudin et al., 2018; Hunt & Reffert, 2023). These achievements have provided new insights into the formation and evolution of stellar clusters (see e.g., Meingast & Alves, 2019; Swiggum et al., 2024), helped constrain the structure and history of the Local Bubble (Zucker et al., 2022; O’Neill et al., 2024), and led to the identification of previously unknown Galactic structures, such as the Radcliffe wave (see e.g., Alves et al., 2020), along with their relation to the past trajectory of the Solar System (see e.g., Maconi et al., 2025).
@@
Concurrent to these astronomical advancements, studies of long-lived radionuclides, like 60Fe (; Wallner et al. 2015a) in geological archives have revealed signatures indicative of nearby SNe or encounters with interstellar regions enriched with these elements (see e.g., Knie et al., 1999; Wallner et al., 2015b, 2021; Koll et al., 2019). In addition, cosmogenic nuclides such as 14C (; Godwin 1962) and 10Be (; Chmeleff et al. 2010), are used for archaeological and geological dating on kyr to Myr timescales. Anomalies in their concentration profiles are significant not only for their physical interpretation but also because, if found in multiple independent archives, they may serve as chronological anchor markers (see e.g., Dee & Pope, 2016).
@@
Recently, Koll et al. (2025) reported on the discovery of a 10Be anomaly in deep ocean crusts of the Central and Northern Pacific during the late Miocene. The origin of this anomaly, dated between 11.5 and 9.0 Myr ago and peaking at 10.1 Myr, remains uncertain, and several scenarios are discussed by the authors. As 10Be is produced by cosmic-ray (CR) spallation in the upper atmosphere (see e.g., Webber & Higbie, 2003), one possibility is that a nearby SN may be responsible for the 10Be excess.
@@
In a recent study, Maconi et al. (2025) showed that around 11.5 Myr ago, at the onset of the 10Be anomaly, the Solar System was exiting the Radcliffe Wave, leaving behind the Orion star formation region, a major site of recent star formation. Given the Sun’s proximity at that time to several massive young clusters, a nearby SN event is a plausible explanation for the observed 10Be anomaly. In this work, we test this hypothesis by integrating the orbits of the Sun and a large sample of clusters back in time over the past 20 Myr and by estimating the probability that a SN occurred within a given distance of the Solar System during the 10Be anomaly.
@@ 2 Data
We primarily used the open cluster catalog by Hunt & Reffert (2023), which was constructed by using the HDBSCAN algorithm (McInnes et al., 2017) on the Gaia DR3 astrometric data (Gaia Collaboration et al., 2023). The catalog contains a total of 7166 star clusters and provides a wide range of parameters, including sky positions, parallaxes, proper motions, extinctions, and stellar membership lists (see Hunt & Reffert, 2023, for details).
We complemented this catalog with four additional clusters not included in Hunt & Reffert (2023), namely CWNU 1028, NGC 1977, OC 0340, and UBC 207. These clusters are part of the Radcliffe Wave and were identified and used in previous studies (Konietzka et al., 2024; Maconi et al., 2025), as detailed in Appendix A. Moreover, we updated the stellar membership of six clusters in the Orion region (ASCC 19, ASCC 20, OCSN 56, OCSN 61, OCSN 65, and Theia 13) with additional members identified using the Significance Mode Analysis (SigMA) clustering algorithm (see Ratzenböck et al., 2023a), applied specifically to the Orion region (A. Rottensteiner, in preparation).
By crossmatching individual cluster members with supplementary radial velocity (RV) surveys, we were able to improve the accuracy of the mean kinematic data of the clusters over the Gaia-only values. For our purposes, we only considered clusters with reliable RV measurements () available for at least three member stars per cluster. After imposing these kinematic quality criteria, our sample comprises of 2725 clusters. Further details on the data curation process are provided in Appendix A.
@@ 3.1 Selection of the clusters of interest
We performed a preliminary orbital integration (see Sect. 3.3) of our initial cluster sample, identifying those that approached the Solar System within a threshold distance of 200 pc over the past 20 Myr. This threshold was chosen to ensure completeness, considering that SNe occurring within 100-150 pc may leave detectable traces on Earth (see e.g., Fry et al., 2015). This step reduced the sample to 278 clusters, about 10% of the initial sample.
@@
For each of these 278 clusters, we performed 1000 orbital integrations, varying their initial positions and velocities by Monte Carlo (MC) sampling the uncertainty distributions. We then selected only the clusters that approached the Solar System within 100 pc during the time span of the 10Be anomaly (9–11.5 Myr ago; see Koll et al., 2025). We adopt 100 pc as a conservative threshold, balancing the range at which SNe may leave traces on Earth with the need for proximity to explain the 10Be anomaly (Koll et al., 2025). We do not explicitly account for the travel time between the SN event and the arrival of CRs at Earth, as this interval is negligible compared to the Myr-scale intervals studied here. This resulted in a subset of 71 clusters.
3.2 Age, mass, and SNe estimation
@@
For the subset of 71 selected clusters identified in Sect. 3.1, we estimated ages, present-day masses, completeness-corrected masses, and the probability of SN events. For the age estimation, we used the Chronos111Chronos code: https://github.com/sebastianratzenboeck/Chronos Python package (see Ratzenböck et al., 2023b), which performs a Bayesian fit of theoretical isochrones to the color magnitude diagram of each cluster. We used the PARSEC isochrones models (Bressan et al., 2012; Nguyen et al., 2022) and assumed solar metallicity () for the stars in the clusters. We limited our fit to to as stars fainter than this threshold have been empirically found to be less well described by isochrone models (see e.g., Ratzenböck et al., 2023b; Rottensteiner & Meingast, 2024).
@@
The present-day cluster masses were estimated by adding up the masses of its stellar members, derived from the corresponding best-fit isochrone. We also computed the initial mass of the clusters corrected for incompleteness, as some star members may be missing due to observational limits, limitation of the clustering algorithm, and stellar evolution (see e.g., Meingast et al., 2021; Hunt & Reffert, 2023; Ratzenböck et al., 2023a): following Meingast et al. (2021), we minimized the difference between the present-day mass function and a Kroupa (2001) initial mass function (IMF). To account for the completeness limits of Gaia data (Gaia Collaboration et al., 2023), we performed the minimization within the mass range . The low-mass end limit of the Kroupa IMF was set to to account for objects below the hydrogen-burning limit, while no high-mass bound was imposed to account for massive stars potentially absent due to Gaia brightness limit or past SNe. The IMF was sampled using the IMF Python code 222IMF code: https://github.com/keflavich/imf.
@@ 3.3 Integration of the orbits
We computed the past orbital trajectories of the clusters and the Solar System using the Galactic dynamics package galpy (Bovy, 2015). galpy enables the numerical integration of orbits over different initial conditions (i.e., Galactocentric distance, height of the Sun above the Galactic disk, velocity of the Sun, and velocity of the Local Standard of Rest; LSR) and various models for the Milky Way potential.
@@ 4 Results and Discussion
Various scenarios are considered by Koll et al. (2025) to explain the 10Be anomaly in the late Miocene. Some involve geological processes that could increase 10Be concentrations in ocean water without altering its atmospheric production rate, while others explore the possibility of an actual enhancement in Galactic CR flux reaching Earth. Among the most promising explanations are the onset and intensification of the Antarctic Circumpolar Current as a terrestrial cause, and either a nearby SN or the compression of the heliosphere by the passage of the Sun through a dense interstellar cloud as astrophysical origins. We refer the reader to their work for a more detailed discussion of all proposed scenarios and focus here solely on the SN hypothesis, further motivated by the fact that around the onset of the 10Be anomaly (11.5 Myr ago), the Solar System was leaving the Orion region of the Radcliffe Wave behind (Maconi et al., 2025), and was therefore in proximity to an active star-forming region where several massive clusters were forming or had just formed.
@@
We find that, out of the 2725 open clusters in our initial sample (see Sect. 2), 19 have a probability greater than 1% of hosting at least one SN within 100 pc of the Solar System and between the onset and the peak of the 10Be anomaly (11.5–10.1 Myr ago), and only four exceed this threshold within 70 pc, namely ASCC 20, OCSN 61, OCSN 56, and CWNU 1057. We estimate the total probability of at least one SN event within 35 pc of the Sun to be around 1%, rising to 5.4% at 40 pc, 14% at 50 pc, and 28% at 70 pc. At 100 pc the total probability across all clusters reaches 68%. We note that none of the considered clusters come closer than 20 pc, thus not reaching the critical distance (8–20 pc) within which a SN could cause an extinction event (see e.g., Gehrels et al., 2003; Thomas & Yelland, 2023). In Fig. 1, we show the distances between the four mentioned clusters and the Solar System over the past 20 Myr, their individual SN probabilities, the total SN probability from all clusters during the relevant interval, and the 10Be concentration profile by Koll et al. (2025). In Table 3, we list the 19 clusters of interest together with their SN probability for different threshold distances. Their properties, including ages and masses, are reported in Table 2, while their heliocentric positions and velocities are given in Table 1.
@@
Among the 19 clusters, ASCC 20 and OCSN 61 (also known as OBP-b), both located in Orion, are the dominant contributors. ASCC 20 is the only cluster with a significant contribution within 70 pc (up to %), while OCSN 61 becomes increasingly relevant beyond 70 pc, reaching 29% at 100 pc. For ASCC 20, which reaches a minimum distance of from the Sun around 11.8 Myr ago and remains within 100 pc throughout the 10Be anomaly, we estimated an age of , consistent with values reported in the literature (see e.g., Kos et al., 2019; Maconi et al., 2025), and derive a present day mass of and an incompleteness corrected initial mass of . For OCSN 61, which remains within 100 pc between the onset and peak of the anomaly but never approaches the Sun closer than 60 pc, we estimated an age of , a present day mass of , and an incompleteness corrected initial mass of . For both of these clusters, we updated the stellar membership using the SigMA algorithm (see Sect. 2 and Appendix A), identifying new members compared to the one listed in Hunt & Reffert (2023). We note that even with an improved member selection, the cluster catalogs are still likely incomplete, suggesting that the estimated number of SN events is conservative due to the possible incompleteness of the cluster catalogs.
@@
To account for systematic uncertainties in the geological age dating, we adopt a conservative estimate of Myr, based on the largest uncertainty reported for one of the samples (Crust-3) used by Koll et al. (2025) and previously analyzed by Wallner et al. (2021). We repeated our analysis with two shifted time windows: [12.0, 10.6] Myr ago and [11.0, 9.6] Myr ago. We find that the total SN probability slightly increases for the older time window and slightly decreases for the younger one, driven primarily by the changing in proximity of ASCC 20 and OCSN 61 (see Fig. 2 and Appendix D). These variations do not affect the result that a nearby SN remains a plausible explanation for the 10Be anomaly.
@@
We note that the 10Be anomaly appears as a broad peak rather than a sharp one, as might be expected from a SN. However, the time interval derived from the crust might be broader than the actual signal, due to diffusion and redistribution processes. Moreover, Koll et al. (2025) consider the SN hypothesis a viable scenario only under certain energetic and geometrical conditions. We consider this plausible given the many uncertainties and physical processes involved in a SN and CR production, including the total SN energy (see e.g., Kasen & Woosley, 2009), the efficiency of energy conversion into CR (see e.g., Blasi, 2011), as well as CR acceleration (see e.g., Caprioli, 2012) and transport (see e.g., Amato & Blasi, 2018). Future advancements in the understanding of these processes will help to better constrain the conditions under which a SN could produce such an anomaly.
@@
The presence of other radionuclides could offer further insights into the origin of the 10Be anomaly and potentially support the nearby SN scenario. For example, 60Fe, produced in SNe, has been detected in geological archives and linked to past SN events (Wallner et al., 2021). So far, there is no indication of a concomitant 60Fe signal with the 10Be anomaly. However, 10Be enhancements have not been observed in association with previously reported 60Fe peaks and this apparent lack of correlation remains an open question. Additionally, the detection of 60Fe within the time window considered here is further complicated by its advanced decay and the expected low concentrations. Among other cosmogenic radionuclides, 53Mn () is a valuable candidate for complementary investigation alongside 10Be, and future accelerator mass spectrometry facilities may enable its detection in the relevant time window (see Koll et al., 2025, for further discussion).
@@
In conclusion, we find that a nearby SN remains a plausible explanation for the 10Be anomaly, especially given the Solar System’s proximity to the Orion region during that period. The estimated SN probability is nonzero at 35 pc and increases with distance, with ASCC 20 and OCSN 61 emerging as the most promising candidate clusters. ASCC 20 is the primary contributor up to 70 pc, while OCSN 61 becomes more relevant beyond that distance. Future investigations of 10Be records from terrestrial archives outside the Pacific Ocean will be crucial to determine whether the observed anomaly reflects a global signal or a regional effect confined to this basin, helping to constrain its terrestrial or astrophysical origin.

### v2

@@
Recent measurements of cosmogenic 10Be in deep-ocean ferromanganese crusts from the central and northern Pacific have revealed an anomalous concentration between 11.5 and 9.0 Myr ago, peaking at 10.1 Myr. One possible explanation is a nearby supernova (SN) event. Motivated by this and by the proximity of the Solar System to the Orion star-forming region during that period, we estimated the probability that at least one SN occurred between the onset and peak of the anomaly. Using an open cluster catalog based on Gaia DR3, we traced back the orbits of 2725 clusters and the Sun over the past 20 Myr and computed the expected number of SN events. We found 19 clusters with a probability greater than 1% each of producing at least one SN within 100 pc of the Sun in the time interval 11.5–10.1 Myr ago. The total cumulative probability exceeds zero at 35 pc from the Sun and increases rapidly with distance, reaching 68% near 100 pc. Two young clusters dominate the SN probability: ASCC 20 contributes most within 70 pc, while OCSN 61 becomes more significant beyond that distance. Our results support the possibility of an SN origin for the 10Be anomaly and highlight the importance of additional 10Be records from independent terrestrial archives to determine whether the anomaly is of astrophysical or terrestrial origin.
@@
The Solar System orbits the center of the Milky Way, together with billions of other stars and vast reservoirs of interstellar gas. Characterizing past environments crossed by the Sun, including possible encounters with large-scale Galactic structures or nearby supernovae (SNe), helps to find possible connections between Galactic environments and Earth’s geological records, fostering interdisciplinary research (e.g., Fuchs et al., 2006; Breitschwerdt et al., 2016; Koll et al., 2019; Wallner et al., 2021; Miller & Fields, 2022; Opher et al., 2024; Maconi et al., 2025; Zucker et al., 2025).
@@
From an astronomical perspective, the European Space Agency’s Gaia mission (Gaia Collaboration et al., 2016) has enabled significant advancements, revolutionizing our understanding of the local Galactic environment. For example, Gaia data have been used to unveil the 3D structure of the solar neighborhood (see e.g., Leike et al., 2020; Vergely et al., 2022; Edenhofer et al., 2024), compile new molecular cloud catalogs (see e.g., Zucker et al., 2019; Cahlon et al., 2024), and significantly expand the open cluster census (see e.g., Castro-Ginard et al., 2018; Cantat-Gaudin et al., 2018; Hunt & Reffert, 2023). These achievements have provided new insights into the formation and evolution of stellar clusters (see e.g., Meingast & Alves, 2019; Swiggum et al., 2024), helped constrain the structure and history of the Local Bubble (Zucker et al., 2022; O’Neill et al., 2024), and led to the identification of previously unknown Galactic structures, such as the Radcliffe wave (Alves et al., 2020), along with their relation to the past trajectory of the Solar System (see e.g., Maconi et al., 2025).
@@
Concurrent to these astronomical advancements, studies of long-lived radionuclides such as 60Fe (; Rugel et al. 2009; Wallner et al. 2015) in geological archives have revealed signatures indicative of nearby SNe or encounters with interstellar regions enriched with these elements (see e.g., Knie et al., 1999; Wallner et al., 2016, 2021; Koll et al., 2019). In addition, cosmogenic nuclides such as 14C (; Kutschera 2013) and 10Be (; Chmeleff et al. 2010; Korschinek et al. 2010), are used for archaeological and geological dating on kiloyear to million-year timescales. Anomalies in their concentration profiles are important both for their physical interpretation and as potential chronological anchor markers if found in multiple independent archives (see e.g., Dee & Pope, 2016).
@@
Recently, Koll et al. (2025) reported on the discovery of a 10Be anomaly in deep ocean crusts of the central and northern Pacific during the late Miocene. The origin of this anomaly, dated between 11.5 and 9.0 Myr ago and peaking at 10.1 Myr, remains uncertain, and several scenarios have been discussed by the authors. As 10Be is produced by cosmic-ray (CR) spallation in the upper atmosphere (see e.g., Webber & Higbie, 2003), one possibility is that a nearby SN may be responsible for the 10Be excess.
@@
In a recent study, Maconi et al. (2025) showed that around 11.5 Myr ago, at the onset of the 10Be anomaly, the Solar System was exiting the Radcliffe wave, leaving behind the Orion star-forming region, where 10-20 SNe likely occurred over the past 12 Myr (see e.g., Bally, 2008). Given the Sun’s proximity at that time to several massive young clusters, a nearby SN event is a possible explanation for the observed 10Be anomaly. In this work, we test this hypothesis by integrating the orbits of the Sun and a large sample of clusters back in time over the past 20 Myr and by estimating the probability that a SN occurred within a given distance of the Solar System during the 10Be anomaly.
@@ 2 Data
We primarily used the open cluster catalog by Hunt & Reffert (2023), which is based on the Gaia DR3 astrometric data (Gaia Collaboration et al., 2023) and contains a total of 7166 star clusters (see Hunt & Reffert, 2023, for details). We complemented this catalog with four additional clusters not included in Hunt & Reffert (2023); namely, CWNU 1028, NGC 1977, OC 0340, and UBC 207. These clusters are part of the Radcliffe wave and were identified and used in previous studies (Konietzka et al., 2024; Maconi et al., 2025), as is detailed in Appendix A. Moreover, we updated the stellar memberships of six clusters in the Orion region (ASCC 19, ASCC 20, OCSN 56, OCSN 61, OCSN 65, and Theia 13) with additional members identified using the significance mode analysis (SigMA) clustering algorithm (Ratzenböck et al., 2023a), applied specifically to the Orion region (A. Rottensteiner, in preparation).
By cross-matching individual cluster members with supplementary radial velocity (RV) surveys, we were able to improve the accuracy of the mean kinematic data of the clusters over the Gaia-only values. For our purposes, we only considered clusters with reliable RV measurements () available for at least three member stars per cluster. After imposing these kinematic quality criteria, our sample comprises 2725 clusters. Further details on the data curation are provided in Appendix A.
@@ 3.1 Selection of the clusters of interest
We performed a preliminary orbital integration (see Sect. 3.3) of our initial cluster sample, identifying those that approached the Solar System within a threshold distance of 200 pc over the past 20 Myr. This threshold was chosen to ensure completeness, considering that SNe occurring within 100-150 pc may leave detectable traces on Earth (see e.g., Fry et al., 2015). This step reduced the sample to 278 clusters (10% of the initial sample).
@@
For each of these 278 clusters, we performed 1000 orbital integrations, varying their initial positions and velocities by Monte Carlo (MC) sampling the uncertainty distributions. We then selected only the clusters that approached the Solar System within 100 pc during the time span of the 10Be anomaly (9–11.5 Myr ago; see Koll et al., 2025). We adopt 100 pc as a conservative threshold, balancing the range at which SNe may leave traces on Earth with the need for proximity to explain the 10Be anomaly (Koll et al., 2025). We do not explicitly account for the travel time between the SN event and the arrival of CRs at Earth, as this interval is negligible compared to the million-year-scale intervals studied here. This resulted in a subset of 71 clusters.
3.2 Age, mass, and SN estimation
@@
For the subset of 71 selected clusters identified in Sect. 3.1, we estimated ages, present-day masses, completeness-corrected masses, and the probability of SN events. For the age estimation, we used the Chronos111Chronos code: https://github.com/sebastianratzenboeck/Chronos Python package (see Ratzenböck et al., 2023b), which performs a Bayesian fit of theoretical isochrones to the color magnitude diagram of each cluster. We used the PARSEC isochrones models (Bressan et al., 2012; Nguyen et al., 2022) and assumed solar metallicity () for the stars in the clusters. We limited our fit to as stars fainter than this threshold have been empirically found to be less well described by isochrone models (see e.g., Ratzenböck et al., 2023b; Rottensteiner & Meingast, 2024).
@@
The present-day cluster masses were estimated by adding up the masses of its stellar members, derived from the corresponding best-fit isochrone. We also computed the initial mass of the clusters corrected for incompleteness, as some star members may be missing due to observational limits, limitation of the clustering algorithm, and stellar evolution (see e.g., Meingast et al., 2021; Hunt & Reffert, 2023; Ratzenböck et al., 2023a). Following Meingast et al. (2021), we minimized the difference between the present-day mass function and a Kroupa (2001) initial mass function (IMF). To account for the completeness limits of Gaia data, we performed the minimization within the mass range . The low-mass end limit of the Kroupa IMF was set to to account for objects below the hydrogen-burning limit, while no high-mass bound was imposed to account for massive stars potentially absent due to Gaia brightness limit or past SNe. The IMF was sampled using the IMF Python code222IMF code: https://github.com/keflavich/imf.
@@ 3.3 Integration of the orbits
We computed the past orbital trajectories of the clusters and the Sun using the Galactic dynamics package galpy (Bovy, 2015). This package enables the numerical orbit integration for different initial conditions (i.e., Galactocentric distance, solar height above the disk, velocity of the Sun, and velocity of the local standard of rest (LSR)) and various models for the Milky Way potential.
@@ 4 Results and discussion
Various scenarios are considered by Koll et al. (2025) to explain the 10Be anomaly in the late Miocene. Some involve geological processes that could increase 10Be concentrations in ocean water without altering its atmospheric production rate, while others explore the possibility of an actual enhancement in Galactic CR flux reaching Earth. Among the most promising explanations are the onset and intensification of the Antarctic circumpolar current as a terrestrial cause, and either a nearby SN or the compression of the heliosphere by the passage of the Sun through a dense interstellar cloud as astrophysical origins. We refer the reader to their work for a more detailed discussion of all proposed scenarios and focus here solely on the SN hypothesis, further motivated by the fact that around the onset of the 10Be anomaly (11.5 Myr ago), the Solar System was leaving the Orion region of the Radcliffe wave behind (Maconi et al., 2025), and was therefore in proximity to an active star-forming region where several massive clusters were forming or had just formed. The difference from Maconi et al. (2025) is that here we consider all high-quality clusters in the solar neighborhood, rather than focusing only on those associated with the Radcliffe wave.
@@
We find that, out of the 2725 open clusters in our initial sample (see Sect. 2), 19 have a probability greater than 1% of hosting at least one SN within 100 pc of the Solar System and between the onset and the peak of the 10Be anomaly (11.5–10.1 Myr ago), and only four exceed this threshold within 70 pc; namely, ASCC 20, OCSN 61, OCSN 56, and CWNU 1057. We estimate the total probability of at least one SN event within 35 pc of the Sun to be around 1%, rising to 5.4% at 40 pc, 14% at 50 pc, and 28% at 70 pc. At 100 pc the total probability across all clusters reaches 68%. We note that none of the considered clusters come closer than 20 pc, and thus none of them reach the critical distance (8–20 pc) within which a SN could cause an extinction event (see e.g., Gehrels et al., 2003; Thomas & Yelland, 2023). In Fig. 1, we show the distances between the four mentioned clusters and the Solar System over the past 20 Myr, their individual SN probabilities, the total SN probability from all clusters during the relevant interval, and the 10Be concentration profile by Koll et al. (2025). In Table 3, we list the 19 clusters of interest together with their SN probability for different threshold distances. Their properties, including ages and masses, are reported in Table 2, while their heliocentric positions and velocities are given in Table 1.
@@
Among the 19 clusters, ASCC 20 and OCSN 61 (also known as OBP-b), both located in Orion, are the dominant contributors. ASCC 20 is the only cluster with a significant contribution within 70 pc (up to %), while OCSN 61 becomes increasingly relevant beyond 70 pc, reaching 29% at 100 pc. For ASCC 20, which reaches a minimum distance of from the Sun around 11.8 Myr ago and remains within 100 pc throughout the 10Be anomaly, we estimated an age of , consistent with values reported in the literature (see e.g., Kos et al., 2019; Maconi et al., 2025), and derive a present day mass of and an incompleteness-corrected initial mass of . For OCSN 61, which remains within 100 pc between the onset and peak of the anomaly but never approaches the Sun closer than 60 pc, we estimated an age of , a present-day mass of , and an incompleteness-corrected initial mass of . For both of these clusters, we updated the stellar membership using the SigMA algorithm (see Sect. 2 and Appendix A), identifying new members compared to the one listed in Hunt & Reffert (2023). Even with improved member selection, cluster catalogs are likely still incomplete, probably making our SN estimates conservative. In Appendix E, we assess the impact of the membership lists on the SN probability estimation.
@@
To account for systematic uncertainties in the geological age dating, we adopt a conservative estimate of Myr, based on the largest uncertainty reported for one of the samples (Crust-3) used by Koll et al. (2025) and previously analyzed by Wallner et al. (2021). We repeated our analysis with two shifted time windows: [12.0, 10.6] Myr ago and [11.0, 9.6] Myr ago. We find that the total SN probability increases slightly in the older time window and decreases in the younger one, mainly due to changes in the proximity of ASCC 20 and OCSN 61 (see Fig. 2 and Appendix D). These variations do not affect the result that a nearby SN remains a possible explanation for the 10Be anomaly.
@@
We note that the 10Be anomaly appears as a broad peak rather than a sharp one, as might be expected from an SN. However, the time interval derived from the crust might be broader than the actual signal, due to diffusion and redistribution processes. Moreover, Koll et al. (2025) consider the SN hypothesis a viable scenario only under certain energetic and geometrical conditions. We consider this possible given the many uncertainties and physical processes involved in a SN and CR production, including the total SN energy (see e.g., Kasen & Woosley, 2009), the efficiency of energy conversion into CR (see e.g., Blasi, 2011), as well as CR acceleration (see e.g., Caprioli, 2012) and transport (see e.g., Amato & Blasi, 2018). Future advancements in the understanding of these processes will help to better constrain the conditions under which an SN could produce such an anomaly.
@@
The presence of other radionuclides could offer further insights into the origin of the 10Be anomaly and potentially support the nearby SN scenario. For example, 60Fe, produced in SNe, has been detected in geological archives and linked to past SN events (Wallner et al., 2021). For the 10Be anomaly analyzed here, no concomitant 60Fe peak has been detected yet. However, 10Be enhancements have not been observed in association with previously reported 60Fe peaks, and this apparent lack of correlation remains an open question. This may be explained by the distances of the SNe from the Solar System; for example, greater than 80 pc for one of the 60Fe anomalies (Breitschwerdt et al., 2016; Schulreich et al., 2023). Additionally, the detection of 60Fe within the time window considered here is further complicated by its advanced decay and the expected low concentrations. Among other cosmogenic radionuclides, 53Mn () is a valuable candidate for complementary investigation alongside 10Be, and future accelerator mass spectrometry facilities may enable its detection in the relevant time window (see e.g., Koll et al., 2025).
@@
In conclusion, we find that a nearby SN remains a possible explanation for the 10Be anomaly, especially given the Solar System’s proximity to the Orion region during that period. The estimated SN probability is nonzero at 35 pc and increases with distance, with ASCC 20 and OCSN 61 emerging as the most promising candidate clusters. ASCC 20 is the primary contributor up to 70 pc, while OCSN 61 becomes more relevant beyond that distance. Future investigations of 10Be records from terrestrial archives outside the Pacific ocean will be crucial to determine whether the observed anomaly reflects a global signal or a regional effect confined to this basin, helping to constrain its terrestrial or astrophysical origin.